Solve by setting up the proper equation to describe the facts given and then carrying out the mathematical calculations to solve for the unknown variable(s).
Three times a larger number added to 4 times a smaller number totals 24. When twice the smaller number is subtracted from twice the larger number, the result is 2. What are the two numbers?

Answer by checkley77(12844)

3x+4y=24
2x-2y=2 or 2x=2y+2 or x=2y/2+2/2 or x=y+1
3(y+1)+4y=24
3y+3+4y=24
7y=24-3
7y=21
y=21/7
y=3 answer.
3x+4*3=24
3x+12=24
3x=24-12
3x=12
x=12/3
x=4 answer.
proof:
2*4-2*3=2
8-6=2
2=2